OpenStudy (anonymous):

What is the y-intercept of the equation y=2(3)-x? 1.x 2.3 3.2 4.-1

OpenStudy (anonymous):

None of it..

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Are you sure your equation is not: y = 2x + 3 ??

OpenStudy (anonymous):

is that

OpenStudy (anonymous):

@best.shakir can you show how you got -1 ??

OpenStudy (anonymous):

it has to be one those are all the answers on the multiple chose the x is a exponet

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=find+y-intercept+of++y%3D2%283%29-x%3F i am not satisfied with this ans

OpenStudy (anonymous):

@best.shakir it is better for you if you will use wolfram alpha for problems harder than this.. You always provide that link.. Try to do it here on your own..

OpenStudy (anonymous):

my steps will not understand to everybody.so that i m sending links

OpenStudy (anonymous):

Yes, please stop citing web resources. I have yet to see an answer that does not involve a direct link to a web resource. If you can arrive at an answer it should be explainable.

OpenStudy (anonymous):

i know that wolfram sometimes will make wrong ans

OpenStudy (anonymous):

@greener6 is it ?? \[\large y = 2(3)^{-x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it has to be 2 right

OpenStudy (anonymous):

So anything raised to the power except 0 is 1.. So: \[y = (2)(3)^0 \implies 2 \times 1 = 2\]

OpenStudy (anonymous):

*Anything raised to the power 0 except 0 is 1..